# Detection of New Galaxy Candidates at 11 in the JADES Field Using JWST NIRCam

### v2

@@ Abstract
We report the detection of seven new galaxy candidates with redshift 11 within the JWST Advanced Deep Extragalactic Survey (JADES) GOODS-S and GOODS-N fields. These new candidates are detected through meticulous analysis of NIRCam photometry in eight filters spanning a wavelength range of 0.8-5.0 m. Photometric redshifts of these galaxy candidates are independently measured utilizing spectral energy distribution (SED) fitting techniques using EAZY and BAGPIPES codes, followed by visual scrutiny. Two of these galaxy candidates are located in GOODS-S field, while the remaining five galaxies are located in GOODS-N field. Our analysis reveals that the stellar masses of these galaxies typically range from log / = 7.75–8.75. Futhermore, these galaxies are typically young with their mass-weighted ages spanning from 80 to 240 Myr. Their specific star formation rates (sSFR) , quantified as ), are measured to vary between to 1.46. These new galaxy candidates offer a robust sample for probing the physical properties of galaxies within the first few hundred Myr of the history of the Universe. We also analyze the relationship between star formation rate (SFR) and stellar mass () within our sample. Using linear regression, our analysis yields a slope of , which we then compare with results from previous studies. Continued investigation through spectroscopic analysis using JWST/NIRSpec is needed to spectroscopically confirm these high-redshift galaxy candidates and investigate further into their physical properties. We plan to follow up on these candidates with future NIRSpec observations.
@@
High-redshift galaxies – JWST1 Introduction
The galaxies that emerged within the first few hundred million years after recombination are believed to have played a crucial role in shaping the early Universe by hosting its first stars. This first generation of stars initiated the process of ionizing neutral hydrogen throughout the cosmos. While reionization is generally understood to have occurred within the first billion years following the Big Bang, the precise mechanisms and the specific types of galaxies responsible for this phenomenon remain largely elusive and subject to ongoing debate (e.g., Barkana & Loeb, 2001; Bromm & Larson, 2004; Fan et al., 2006; Bromm & Yoshida, 2011; Stark, 2016; Ouchi et al., 2020). Thus, studying early galaxies is crucial for understanding their formation and evolution over billions of years, revealing insights into the physical processes that shaped them and their role in re-ionizing the Universe.
@@
The method for selecting high-redshift galaxies has primarily been driven by photometry (Schenker et al., 2013; Bouwens et al., 2015). This technique exploits the absorption of ultraviolet light by neutral hydrogen within, around, and between distant galaxies. This absorption produces a distinctive spectral feature known as the “Lyman break," observable at 912 Å in the spectral energy distribution (SED) of the galaxy (Sarkar & Samui, 2019). The redshifts of these galaxies are determined by fitting their observed SEDs to either simulated or observed galaxy SEDs with known physical properties. This method incorporates more data than pure color selection alone (Bolzonella et al., 2000). To effectively apply this approach, accurate template SEDs covering the entire color space of photometry are necessary, accounting for factors such as dust extinction and intergalactic medium (IGM) absorption. However, this procedure becomes less certain for high-redshift galaxies due to the scarcity of ultraviolet (UV) and optical SEDs for early galaxies (e.g., Stark, 2016).
@@
The commissioning of the James Webb Space Telescope (JWST) (Rigby et al., 2023) has opened a new window into the discovery and study of galaxies during the earliest stages of the universe, made possible by its improved sensitivity, 7 light-collecting area, and the broad imaging field of NIRCam (Rieke et al., 2005). In the early months of operation, several 11 galaxy candidates were identified (Naidu et al., 2022; Finkelstein et al., 2022; Harikane et al., 2023; Donnan et al., 2023; Atek et al., 2023a) as part of JWST’s Early Release Observations (ERO; Pontoppidan et al. (2022)), the Early Release Science (ERS) programs CEERS (Bagley et al., 2023) and Through the Looking GLASS (GLASS-JWST) programs (Treu et al., 2022). Many of the galaxy candidates identified in these early observations exceeded the previous distance record established by HST studies (Oesch et al., 2016). The discovery of early galaxy candidates has been further advanced by observational campaigns such as the JWST Advanced Deep Extragalactic Survey (JADES, Eisenstein et al., 2023a) , the Next Generation Deep Extragalactic Exploratory Public (NGDEEP, Austin et al., 2023) survey, and the Prime Extragalactic Areas for Reionization and Lensing Science (PEARLS) project (Windhorst et al., 2023; Juodžbalis et al., 2023), which have utilized photometric techniques to identify z 9-16 galaxy candidates.
@@
JADES is one of the most extensive initiatives in JWST’s first year of observation, dedicating 32 days of telescope time to the discovery and characterization of the most distant galaxies (Rieke et al., 2023a). The survey encompasses two deep fields, GOODS-South and GOODS-North, which have extensive legacy datasets provided by the Great Observatories Origins Deep Survey (GOODS; Dickinson et al., 2003; Giavalisco et al., 2004), and utilizes three instruments: NIRCam for imaging, NIRSpec for spectroscopy, and MIRI with an imaging and spectroscopy. JADES operates at two levels of survey depth: "Deep" and "Medium." The Deep survey is designed for in-depth characterization of fainter galaxies or detailed analysis of individual sources at higher signal-to-noise ratios, whereas the Medium survey facilitates a statistical analysis of the high-redshift galaxy population.
@@
In this paper, we report the detection of new 11 galaxy candidates using NIRCam observations taken across eight filters from the first year of the JADES survey in the GOODS-S and GOODS-N fields. We use two distinct codes EAZY (Brammer et al., 2008) and BAGPIPES (Carnall et al., 2018) for determining the redshifts of the candidates and their physical properties using SED fitting techniques to the photometric data. The format of the paper is as follows. In Section 2, we provide a detailed description of the observation and the photometric data reduction process, including photometric redshift estimation and dropout selection methods. In section 3, we detail the methodology for selecting galaxies at redshifts 11. In Section 4, we present the main results from our galaxy sample, including detailed descriptions, sizes, properties derived from SED fitting, and an analysis of systematic uncertainties and potential sources of contamination. In Section 5, we summarize our findings.
@@
Throughout this paper, we adhere to the AB magnitude system (Oke & Gunn, 1983) and cosmological parameters reported by Planck Collaboration et al. (2020): the Hubble constant , matter density parameter , and dark energy density .
@@ 2 Observation and Methods
The debut of the JWST in December 2021 opens up an unprecedented window into the study of the distant cosmos. Throughout the past year, a series of observation initiatives have been underway e.g., JADES (Hainline et al., 2024), CHEERS (Finkelstein et al., 2015), GLASS (Castellano et al., 2022), NGDEEP (Leung et al., 2023) and many others. This paper centers on the analysis of data stemming from the JADES programs, with a particular focus on photometry.
@@ 2.1 JADES GOODS-S and GOODS-N observations
In this study, we investigate galaxy candidates identified through NIRCam imaging within the GOODS-S and GOODS-N field. GOODS-S encompasses observations conducted under Program 1180 led by PI Eisenstein from UT 2022-09-29 to 2022-10-10 (e.g., Bunker et al., 2023; Eisenstein et al., 2023a; Rieke et al., 2023b; Eisenstein et al., 2023b). Additionally, NIRCam parallels, covering 9.8 square arcminutes each, were observed during NIRSpec PID 1210 led by PI Ferruit from UT 2022-10-20 to 2022-10-24, positioned within and southwest of the JADES Medium footprint in GOODS-S. Further NIRCam observations (9.8 square arcminutes) parallel to NIRSpec PID 1286, also led by PI Ferruit, occurred from UT 2023-01-12 to 2023-01-13, situated northwest of the JADES Deep footprint in GOODS-S (e.g., Robertson et al., 2023; Tacchella et al., 2023; Hainline et al., 2024). The total survey area for JADES GOODS-S spans 67 square arcminutes, with 27 square arcminutes allocated for the JADES Deep program and 40 square arcminutes for the JADES Medium program. Utilizing filters F090W, F115W, F150W, F200W, F277W, F335M, F356W, F410M, and F444W ( = 0.8 – 5.0m) for JADES Deep and the same filters without F335M for JADES Medium, these observations incorporate a minimum of six dither points per observation, with exposure times ranging from 14 to 60 ks. Consequently, the 5 depths range from 3.4 to 5.9 nJy, with flux aperture sizes varying between 1.26 and 1.52 arcseconds.
@@
The NIRCam GOODS-N program encompasses a total area of 58 square arcminutes. It’s split into two sections: the northwest (NW) covering 30.4 square arcminutes, and the southeast (SE) covering 27.6 square arcminutes. The NW part was observed as part of PID 1181 (PI: Eisenstein), using NIRCam as the main instrument. JADES guarantees a high level of pixel diversity across all filter bands, effectively mitigating flat-field inaccuracies, cosmic-ray interference, and other pixel-level issues.
@@
For this study, we utilized publicly available JADES data and reductions111https://archive.stsci.edu/hlsp/jades. We specifically used NIRCam images from eight filters: F090W, F115W, F150W, F200W, F277W, F356W, F410M, and F444W. These images were processed with a pixel scale of 0.03 per pixel. To identify sources in the images, we created a detection image by combining the F200W and F277W NIRCam bands using a weighted sum. We used image-segmentation tools from Photutils (Bradley et al., 2023a) and JWST data analysis tool notebooks 222https://jwst-docs.stsci.edu/jwst-post-pipeline-data-analysis to identify sources with a SNR above 3 across five contiguous pixels. From this source catalog, we measured elliptical Kron aperture fluxes using Photutils. We used a flexible Kron scale factor (K = 1.5) and a circularized radius six times larger than the Gaussian-equivalent elliptical sizes to ensure accurate measurements while masking any segmentation regions of neighboring sources (Bradley et al., 2023b; Hainline et al., 2024). The size of elliptical Kron aperture for each source was estimated by multiplying the Kron scale factor by the individual Kron radius.
@@
We extracted and fitted to SED for each source adopting the Kron fluxes, as previous studies have demonstrated that measuring colors in small elliptical apertures accurately captures the colors of distant galaxies (Finkelstein et al., 2023; Bradley et al., 2023b). Additionally, we repeated a similar analysis using Kron apertures with a scale factor of 2.5. To correct for aperture effects, we estimated an aperture correction by comparing the total fluxes measured in the larger and smaller apertures for each source. This correction factor was then applied to the fluxes and uncertainties for all filters.
@@ 2.2 Photometric redshift estimation
We used EAZY-py, the Python implementation of the EAZY photometric redshift estimator code (Brammer et al., 2008), to derive the initial photometric redshifts. We utilize a set of 12 “tweakfspsQSF12v3 templates derived from the Flexible Stellar Population Synthesis (FSPS) library (Conroy et al., 2009; Conroy & Gunn, 2010). These templates include a range of galaxy types including star-forming, quiescent, and dusty and also include a range of star-formation histories, e.g., bursty, slowly rising, and slowly falling. Additionally, we utilize six newly developed templates from Larson et al. (2023), which were also applied by Adams et al. (2024) for deriving photometric redshifts of galaxies with z > 7.5. These additional templates were found to provide better photometric redshift estimation for 9 galaxies.
We utilized the values generated by EAZY to compute a probability distribution, P(), assuming a uniform redshift prior. This distribution is given by the equation , and it is normalized such that the integral of P() over all redshifts equals 1.0. We vary redshifts in 0.1 20, in steps of 0.01. Since the high-redshift luminosity funciton is still not well-constrained, we adpoted a flat luminosity prior to prevent bias against the selection of bright high-redshift galaxies. Similar techniques were also adopted in several past JWST studies, such as Finkelstein et al. (2022); Bradley et al. (2023b).
3 Selection criteria for high-redshift candidate
We selected high-redshift galaxy candidates by combining EAZY-based SED fitting for photometric redshifts, signal-to-noise (SNR) criteria, as adopted in several past studies such as Salmon et al. (2020); Finkelstein et al. (2022, 2023); Bradley et al. (2023b); Leung et al. (2023); Atek et al. (2023a); Donnan et al. (2023); Hainline et al. (2024). We adopted photometric SNR criteria to ensure nondetections in filters bluer than Ly and robust detections in filters redder than Ly break. Finally, we visually inspect each galaxy candidates for detector artifacts, diffraction spikes, and spurious noise close to detector edge.
Our selection criteria for identifying galaxy candidates with 11 are similar to the criteria adopted in Finkelstein et al. (2022) and Bradley et al. (2023b)-

### v3

@@ Abstract
We report the detection of six new galaxy candidates with photometric redshifts 11 within the JWST Advanced Deep Extragalactic Survey (JADES) GOODS-S and GOODS-N fields. These new candidates are detected through meticulous analysis of NIRCam photometry in eight filters spanning a wavelength range of 0.8-5.0 m. Photometric redshifts of these galaxy candidates are independently measured utilizing spectral energy distribution (SED) fitting techniques using EAZY and BAGPIPES codes, followed by visual scrutiny. One of these galaxy candidates is located in GOODS-S field, while the remaining five galaxies are located in GOODS-N field. Our analysis reveals that the stellar masses of these galaxies typically range from log / = 7.75–8.75. Futhermore, these galaxies are typically young with their mass-weighted ages spanning from 80 to 240 Myr. Their specific star formation rates (sSFR) , quantified as ), are measured to vary between to 1.46. These new galaxy candidates offer a robust sample for probing the physical properties of galaxies within the first few hundred Myr of the history of the Universe. We also analyze the relationship between star formation rate (SFR) and stellar mass () within our sample. Continued investigation through spectroscopic analysis using JWST/NIRSpec is needed to spectroscopically confirm these high-redshift galaxy candidates and investigate further into their physical properties. We plan to follow up on these candidates with future NIRSpec observations.
@@
galaxies: high-redshift, galaxies: photometry, galaxies: distances and redshifts, infrared: galaxies1 Introduction
The galaxies that emerged within the first few hundred million years after recombination are believed to have played a crucial role in shaping the early Universe by hosting its first stars. This first generation of stars initiated the process of ionizing neutral hydrogen throughout the cosmos. While reionization is generally understood to have occurred within the first billion years following the Big Bang, the precise mechanisms and the specific types of galaxies responsible for this phenomenon remain largely elusive and subject to ongoing debate (e.g., Barkana and Loeb, 2001; Bromm and Larson, 2004; Fan et al., 2006; Bromm and Yoshida, 2011; Stark, 2016; Ouchi et al., 2020). Thus, studying early galaxies is crucial for understanding their formation and evolution over billions of years, revealing insights into the physical processes that shaped them and their role in re-ionizing the Universe.
@@
The method for selecting high-redshift galaxies has primarily been driven by photometry (Schenker et al., 2013; Bouwens et al., 2015). This technique exploits the absorption of ultraviolet light by neutral hydrogen within, around, and between distant galaxies. This absorption produces a distinctive spectral feature known as the “ Lyman- break," observable at 912 Å in the spectral energy distribution (SED) of the galaxy (Sarkar and Samui, 2019). The redshifts of these galaxies are determined by fitting their observed SEDs to either simulated or observed galaxy SEDs with known physical properties. This method incorporates more data than pure color selection alone (Bolzonella et al., 2000). To effectively apply this approach, accurate template SEDs covering the entire color space of photometry are necessary, accounting for factors such as dust extinction and intergalactic medium (IGM) absorption. However, this procedure becomes less certain for high-redshift galaxies due to the scarcity of ultraviolet (UV) and optical SEDs for early galaxies (e.g., Stark, 2016).
@@
The commissioning of the James Webb Space Telescope (JWST) (Rigby et al., 2023) has opened a new window into the discovery and study of galaxies during the earliest stages of the universe, made possible by its improved sensitivity, 7 light-collecting area, and the broad imaging field of NIRCam (Rieke et al., 2005). In the early months of operation, several 11 galaxy candidates were identified (Naidu et al., 2022; Finkelstein et al., 2022; Harikane et al., 2023; Donnan et al., 2023; Atek et al., 2023b) as part of JWST’s Early Release Observations (ERO; Pontoppidan et al. (2022)), the Early Release Science (ERS) programs CEERS (Bagley et al., 2023) and Through the Looking GLASS (GLASS-JWST) programs (Treu et al., 2022). Many of the galaxy candidates identified in these early observations exceeded the previous distance record established by HST studies (Oesch et al., 2016). The discovery of early galaxy candidates has been further advanced by observational campaigns such as the JWST Advanced Deep Extragalactic Survey (JADES, Eisenstein et al., 2023b) , the Next Generation Deep Extragalactic Exploratory Public (NGDEEP, Austin et al., 2023) survey, and the Prime Extragalactic Areas for Reionization and Lensing Science (PEARLS) project (Windhorst et al., 2023; Juodžbalis et al., 2023), which have utilized photometric techniques to identify z 9-16 galaxy candidates.
@@
JADES is one of the most extensive initiatives in JWST’s first year of observation, dedicating 32 days of telescope time to the discovery and characterization of the most distant galaxies (Rieke et al., 2023a). The survey encompasses two deep fields, GOODS-South and GOODS-North, which have extensive legacy datasets provided by the Great Observatories Origins Deep Survey (GOODS; Dickinson et al., 2003; Giavalisco et al., 2004), and utilizes three instruments: NIRCam for imaging, NIRSpec for spectroscopy, and MIRI with an imaging and spectroscopy. JADES operates at two levels of survey depth: "Deep" and "Medium." The Deep survey is designed for in-depth characterization of fainter galaxies or detailed analysis of individual sources at higher signal-to-noise ratios, whereas the Medium survey facilitates a statistical analysis of the high-redshift galaxy population. Recently, Hainline et al. (2024) presents a impressive catalog of 717 candidate galaxies at redshifts , identified using deep imaging from JWST JADES program along with complementary data from FRESCO, JEMS, and HST. The candidates, selected through photometric redshifts, include many newly discovered sources—some potentially at redshifts up to .
@@
In this paper, we report the detection of new 11 galaxy candidates using NIRCam observations taken across eight filters from the first year of the JADES survey in the GOODS-S and GOODS-N fields. The targets analyzed in this study are also included in the catalog released by JADES team, although with lower redshift estimates. Our analysis provides independent redshift measurements for these sources, based on refined photometry and an alternative modeling approach, offering a complementary perspective on their physical properties. We use two distinct codes EAZY (Brammer et al., 2008) and BAGPIPES (Carnall et al., 2018) for determining the redshifts of the candidates and their physical properties using SED fitting techniques to the photometric data. The format of the paper is as follows. In Section 2, we provide a detailed description of the observation and the photometric data reduction process, including photometric redshift estimation and dropout selection methods. In section 3, we detail the methodology for selecting galaxies at redshifts 11. In Section 4, we present the main results from our galaxy sample, including detailed descriptions, sizes, properties derived from SED fitting, and an analysis of systematic uncertainties and potential sources of contamination. In Section 5, we summarize our findings.
@@
Throughout this paper, we adhere to the AB magnitude system (Oke and Gunn, 1983) and cosmological parameters reported by Planck Collaboration et al. (2020): the Hubble constant , matter density parameter , and dark energy density .
@@ 2 Observation and Methods
The debut of the JWST in December 2021 opens up an unprecedented window into the study of the distant cosmos. Throughout the past year, a series of observation initiatives have been underway e.g., JADES (Hainline et al., 2024), CEERS (Finkelstein et al., 2015), GLASS (Castellano et al., 2022), NGDEEP (Leung et al., 2023) and many others. This paper centers on the analysis of data stemming from the JADES programs, with a particular focus on photometry.
@@ 2.1 JADES GOODS-S and GOODS-N observations
In this study, we investigate galaxy candidates identified through NIRCam imaging within the GOODS-S and GOODS-N field. GOODS-S encompasses observations conducted under Program 1180 led by PI Eisenstein from UT 2022-09-29 to 2022-10-10 (e.g., Bunker et al., 2023; Eisenstein et al., 2023b; Rieke et al., 2023b; Eisenstein et al., 2023a). Additionally, NIRCam parallels, covering 9.8 square arcminutes each, were observed during NIRSpec PID 1210 led by PI Ferruit from UT 2022-10-20 to 2022-10-24, positioned within and southwest of the JADES Medium footprint in GOODS-S. Further NIRCam observations (9.8 square arcminutes) parallel to NIRSpec PID 1286, also led by PI Ferruit, occurred from UT 2023-01-12 to 2023-01-13, situated northwest of the JADES Deep footprint in GOODS-S (e.g., Robertson et al., 2023; Tacchella et al., 2023; Hainline et al., 2024). The total survey area for JADES GOODS-S spans 67 square arcminutes, with 27 square arcminutes allocated for the JADES Deep program and 40 square arcminutes for the JADES Medium program. Utilizing filters F090W, F115W, F150W, F200W, F277W, F335M, F356W, F410M, and F444W ( = 0.8 – 5.0m) for JADES Deep and the same filters without F335M for JADES Medium, these observations incorporate a minimum of six dither points per observation, with exposure times ranging from 14 to 60 ks. Consequently, the 5 depths range from 3.4 to 5.9 nJy, with flux aperture sizes varying between 1.26 and 1.52 arcseconds.
@@
The NIRCam GOODS-N program encompasses a total area of 58 square arcminutes. It’s split into two sections: the northwest (NW) covering 30.4 square arcminutes, and the southeast (SE) covering 27.6 square arcminutes. The NW part was observed as part of PID 1181 (PI: Eisenstein), using NIRCam as the main instrument. JADES guarantees a high level of pixel diversity across all filter bands, effectively mitigating flat-field inaccuracies, cosmic-ray interference, and other pixel-level issues. Figure 1 illustrates the GOODS-S and GOODS-N field covered by JADES survey, adopted from Hainline et al. (2024).
@@
For this study, we utilized publicly available JADES data and reductions111https://archive.stsci.edu/hlsp/jades. We specifically used NIRCam images from eight filters: F090W, F115W, F150W, F200W, F277W, F356W, F410M, and F444W. These images were processed with a pixel scale of 0.03 per pixel. To identify sources in the images, we created a detection image by combining the F200W and F277W NIRCam bands using a weighted sum. We used image-segmentation tools from Photutils (Bradley et al., 2023b) and JWST data analysis tool notebooks 222https://jwst-docs.stsci.edu/jwst-post-pipeline-data-analysis to identify sources with a SNR above 3 across five contiguous pixels. From this source catalog, we measured elliptical Kron aperture fluxes using Photutils.
We used a flexible Kron scale factor (K = 1.5) and a circularized radius six times larger than the Gaussian-equivalent elliptical sizes to ensure accurate measurements while masking any segmentation regions of neighboring sources (Bradley et al., 2023a; Hainline et al., 2024). The size of elliptical Kron aperture for each source was estimated by multiplying the Kron scale factor by the individual Kron radius.
@@
We extracted and fitted to SED for each source adopting the Kron fluxes, as previous studies have demonstrated that measuring colors in small elliptical apertures accurately captures the colors of distant galaxies (Finkelstein et al., 2023; Bradley et al., 2023a). Additionally, we repeated a similar analysis using Kron apertures with a scale factor of 2.5. To correct for aperture effects, we estimated an aperture correction by comparing the total fluxes measured in the larger and smaller apertures for each source. This correction factor was then applied to the fluxes and uncertainties for all filters.
@@ 2.2 Photometric redshift estimation
We used EAZY-py, the Python implementation of the EAZY photometric redshift estimator code (Brammer et al., 2008), to derive the initial photometric redshifts. We utilize a set of 12 “tweakfspsQSF12v3” templates derived from the Flexible Stellar Population Synthesis (FSPS) library (Conroy et al., 2009; Conroy and Gunn, 2010). These templates include a range of galaxy types including star-forming, quiescent, and dusty and also include a range of star-formation histories, e.g., bursty, slowly rising, and slowly falling. Additionally, we utilize six newly developed templates from Larson et al. (2023), which were also applied by Adams et al. (2024) for deriving photometric redshifts of galaxies with z > 7.5. These additional templates were found to provide better photometric redshift estimation for 9 galaxies.
We utilized the values generated by EAZY to compute a probability distribution, P(), assuming a uniform redshift prior. This distribution is given by the equation , and it is normalized such that the integral of P() over all redshifts equals 1.0. We vary redshifts in 0.1 20, in steps of 0.01. Since the high-redshift luminosity funciton is still not well-constrained, we adpoted a flat luminosity prior to prevent bias against the selection of bright high-redshift galaxies. Similar techniques were also adopted in several past JWST studies, such as Finkelstein et al. (2022); Bradley et al. (2023a).
3 Selection criteria for high-redshift candidates
We selected high-redshift galaxy candidates by combining EAZY-based SED fitting for photometric redshifts, signal-to-noise (SNR) criteria, as adopted in several past studies such as Salmon et al. (2020); Finkelstein et al. (2022, 2023); Bradley et al. (2023a); Leung et al. (2023); Atek et al. (2023b); Donnan et al. (2023); Hainline et al. (2024). We adopted photometric SNR criteria to ensure nondetections in filters bluer than Ly and robust detections in filters redder than Ly break. Finally, we visually inspect each galaxy candidates for detector artifacts, diffraction spikes, and spurious noise close to detector edge.
Our selection criteria for identifying galaxy candidates with 11 are similar to the criteria adopted in Finkelstein et al. (2022) and Bradley et al. (2023a)-